OpenStudy (anonymous):

Explain how to convert 9.9x10^6 from scientific notation to standard form, then give the answer in standard form.

hero (hero):

You are given \(9.9 \times 10^6\) Also note that: \(10^6 = 1000000\) Replace \(10^6\) with \(1000000\) \(9.9 \times 1000000\) Then multiply: \(9,900,000\)